OpenStudy (anonymous):

Find "x" in 3 sin(x)-sqrt(3) cos(x) = sqrt(3) Please show working out. Thanks! Find "x" in 3 sin(x)-sqrt(3) cos(x) = sqrt(3) Please show working out. Thanks! @Mathematics

OpenStudy (perl):

you have to use a calculator, its a non linear equation ,

OpenStudy (perl):

3sin(x) - sqrt(3)cosx - sqrt(3) = 0

OpenStudy (anonymous):

But we're not allowed to use calculators to solve this.

OpenStudy (perl):

3sin x = sqrt 3 + sqrt 3 cos x

OpenStudy (perl):

3 sin x = sqrt 3 ( 1 + cos x )

OpenStudy (perl):

sqrt 3 sin x= 1 + cos x

OpenStudy (anonymous):

I think we have to solve it by R-method (Auxiliary Angle Method) but it doesn't work when I try it.

OpenStudy (perl):

whats R method?

OpenStudy (anonymous):

umm... do you know sum of products of trigonometric functions?

OpenStudy (perl):

ummm.. yeah i invented them

OpenStudy (anonymous):

lol well it uses that and then somehow you let 3 equal to cos(u) and sin(u) equal to sqrt3... wait look on this site. I can't explain it properly http://trevorpythag.co.uk/2009/mathematics/trigonometry/auxiliary-angle-method-for-solving-trigonometry-equations/

OpenStudy (perl):

ok im reading this , i invented this so its not going to be hard to remember

OpenStudy (anonymous):

ok. =D

OpenStudy (mimi_x3):

asinx - bcosx = rsin(x-a) Then find r = sqrt(a^2+b^2)

OpenStudy (anonymous):

yup I did that. It equals to 2sqrt(3)

OpenStudy (mimi_x3):

then find tan= a/b

OpenStudy (mimi_x3):

tan(a) = a/b*

OpenStudy (anonymous):

1/sqrt(3) ?

OpenStudy (mimi_x3):

3/sqrt(3) i think

OpenStudy (anonymous):

ok ... but wouldn't we make tan(x) by doing b/a?

OpenStudy (perl):

ok we will rewrite the left side of the equation 3 sin(x)-sqrt(3) cos(x) =r sin(x+y) 3 sin(x)-sqrt(3) cos(x) = r sinx cosy + r siny cos x so rcos y = 3, r sin y = -sqrt 3

OpenStudy (mimi_x3):

woops tan(x) = b/a -> sqrt(3)/3 ->30

OpenStudy (anonymous):

@perl why does rsiny=-sqrt3? Why is sqrt3 negative?

OpenStudy (perl):

using the identity, cos^2 + sin^2 = 1, we have [ r cos y ] ^2 + [ r sin y ] ^2 = r^2, 3^2 + (-sqrt 3)^2 = r^2

OpenStudy (perl):

because we are matching the left and right terms

OpenStudy (anonymous):

?

OpenStudy (perl):

if it is true that 3 sin(x)-sqrt(3) cos(x) = r sinx cosy + r siny cos x then it must be true that 3 sin x = r sin x cos y -sqrt 3 cos x = rsin y cos x

OpenStudy (alfie):

There's a easy way to do this. You just substitute: \[sinx = \frac{2t}{1+t^2} , cosx = \frac{1-t^2}{1+t^2}\] Getting = \[6t-\sqrt3 + \sqrt3t^2 = \sqrt3 +\sqrt3t^2 \rightarrow 6t = 2\sqrt3 \rightarrow t = \frac{\sqrt3}{3} \rightarrow \tan(x/2) = \frac{\sqrt3}{3} \rightarrow \] x = \[1/3 (π+6 n π)\]

OpenStudy (perl):

if it is true that 3 sin(x)-sqrt(3) cos(x) = r sinx cosy + r siny cos x then it must be true that 3 sin x = r cos y * sin x -sqrt 3 cos x = rsin y cos x

OpenStudy (alfie):

and π/3+2 n π

OpenStudy (perl):

its like equating coefficients ,

OpenStudy (perl):

alfie, where did you get that identity?

OpenStudy (anonymous):

@alfie Thanks! but that's using T-method. I have to solve this question using R-method

OpenStudy (perl):

or that substitution?

OpenStudy (anonymous):

@perl ok I get why now!

OpenStudy (perl):

hmmm, thats a weird substitution. its true that (2t/(1+t^2)) + ( 1-t^2 /1+t^2)^2 = 1 , but ...

OpenStudy (perl):

duplicity, now rcosy = 3 , and r sin y = -sqrt 3 ,

OpenStudy (anonymous):

@ alfie, what would x equal at the end if it was in degrees?

OpenStudy (anonymous):

@perl ok continue on =D

OpenStudy (agreene):

you do realise you could solve this by substitution, right?

OpenStudy (perl):

and we know using the identity, cos^2 + sin^2 = 1, we have [ r cos y ] ^2 + [ r sin y ] ^2 = r^2, 3^2 + (-sqrt 3)^2 = r^2 , so r = sqrt 12

OpenStudy (perl):

agree, what kind of substitution?

OpenStudy (anonymous):

@agreene yes, but the question said to answer using R-method

OpenStudy (anonymous):

@perl yup r=2sqrt3

OpenStudy (mimi_x3):

r= sqrt(12) a = 30 Therefore, sqrt(12(x+3)= sqrt(3) sin(x+30) = sqrt(3)/sqrt(12) r.angle = sin^-1=(sqrt(3)/sqrt(12) = 30 x + 30 = 180 + 30 , 360-30 x + 30 = 210 , 330 x =...

OpenStudy (perl):

now r cos y = 3, so 2sqrt 3 cos y = 3 , so cos y = 3/ ( 2 sqrt 3 ) = sqrt 3 / 2

OpenStudy (perl):

so y = pi/6 + 2pi*n, 11pi/6 + 2pi*n

OpenStudy (anonymous):

@ perl what's 'n'?

OpenStudy (perl):

any integer, { ...-3,-2,-1,0, 1 , 2 ,...}

OpenStudy (anonymous):

@mimi, Thanks! though... it's hard to read TT_TT

OpenStudy (agreene):

from: 3sin(x)-sqrt(3)cos(x)-sqrt(3)=0 Let u=tan(x/2) sin(x) = 2u/(u^2+1) cos(x)= (1-u^2)/(u^2+1) then we are at: (6u)/(u^2+1)-(sqrt(3)(1-u^2))/(u^2-1)-sqrt(3)=0 do some rearranging: \[-\frac{2(\sqrt{3}-3u)}{u^2-1}\] solving that we get to \[u=\frac{1}{\sqrt{3}}\] sub that kid: \[\tan\frac{x}{2}=\frac{1}{\sqrt{3}}\] inverse tan both sides then multiply by 2 \[\large x=\frac{\pi}{6}\]

OpenStudy (anonymous):

eep! so many answers... so x = 30, 300, 180?

OpenStudy (mimi_x3):

lols yeah they left it in radian form. Not sure if my one is right though, it has been a while that i have done it.

OpenStudy (agreene):

my answer is in radians, so 30 degrees, btw.

OpenStudy (agreene):

but it isnt entirely complete, now that I think of it... there should be infinite solutions... so yeah

OpenStudy (mimi_x3):

is my one one right ? >_<

OpenStudy (mimi_x3):

my one*

OpenStudy (anonymous):

hmm.... o...k... I don't know!!

OpenStudy (anonymous):

oh and @mimi isn't it sin(x-30) not sin(x+30)

OpenStudy (alfie):

solutions cannot be just one angle, they have to be infinite since the tangent is a periodic function, you will have to plug in the period, otherwise that's not correct, or partially correct ;)

OpenStudy (perl):

something is wrong with my math, i got now r cos y = 3, so 2sqrt 3 cos y = 3 , so cos y = 3/ ( 2 sqrt 3 ) = sqrt 3 / 2 y = pi/6 + 2pi*n, 11pi/6 + 2pi*n

OpenStudy (agreene):

\[x=2\pi n+\pi\] And\[n \in\mathbb{Z}\] should be your answer in radians (from my above solution)

OpenStudy (perl):

but if you do rcos y = 3, r sin y = -sqrt 3, then tan y = -sqrt 3 / 3 , that gives me

OpenStudy (perl):

i dont get pi/6 as a solution

OpenStudy (alfie):

@Agreene I agree with you, I did your same reasoning and got the same solutions up there ;)

OpenStudy (anonymous):

hmm... some how I got tan y = 1/sqrt3

OpenStudy (perl):

if i do tan y = -sqrt 3 /3 , the solutions are 5pi/6 + pi*n, but if i do r cos y = 3, so 2sqrt 3 cos y = 3 , so cos y = 3/ ( 2 sqrt 3 ) = sqrt 3 / 2 y = pi/6 + 2pi*n, 11pi/6 + 2pi*n , i get a different answer

OpenStudy (agreene):

@alfie, I musta missed it in my awesome scrolling and wondering what everyone was doing O.o

OpenStudy (anonymous):

eep! I'm confused

OpenStudy (perl):

im getting different solutions for y

OpenStudy (perl):

ok suppose you are given r cos y = 3, r sin y = -sqrt 3 , r = 2 sqrt 3 , you can divide rsin y by r cos y , and get tan y = -sqrt 3/ 3 , or you can solve r cos y = 3 seperately

OpenStudy (perl):

im going to have to post this as a new question, i found a hole in mathematics

OpenStudy (anonymous):

@ perl Wholey!

OpenStudy (perl):

one sec im going to research this

OpenStudy (perl):

mimi, can you help ?

OpenStudy (perl):

suppose I am given sin y = -1/2 and cos y = sqrt(3)/2, I get two different solutions. if i do siny / cos y = tan y = (-1/2)/ (sqrt(3)/2) tan y = -sqrt 3 / 3 , i get 5pi/6 + pi*n. But if i do cos y = sqrt 3 / 2 , i get a different answer

OpenStudy (mimi_x3):

lols, i have no idea when you're doing. You're just making it more complicated.

OpenStudy (perl):

i need to straighten out something first

OpenStudy (perl):

its a side problem

OpenStudy (mimi_x3):

I don't think that the question is that complicated :/

OpenStudy (perl):

mim, its just a different question entirely. suppose you want to solve the system of equations sin y = -1/2 cos y = sqrt(3)/2 , it would be wrong to do tan y = -sqrt 3 / 3 , that comes out with wrong solutions

OpenStudy (mimi_x3):

woudn't it be -2/2sqrt(3) ?

OpenStudy (perl):

(-1/2) / ( sqrt(3)/2) = -sqrt (3)/3

OpenStudy (perl):